### v1

@@ 1 Introduction
Language modeling (LM) is a fundamental approach for achieving cognitive intelligence in the field of natural language processing (NLP), and its progress has been notable in recent years [1; 2]. It assumes a central role in understanding, generating, and manipulating human language, serving as the cornerstone for a diverse range of NLP applications, including machine translation, chatbots, sentiment analysis, and text summarization. With the evolution of deep learning, the early statistical language models (SLM) have gradually transformed into neural language models (NLM) based on neural networks. This shift is characterized by the adoption of word embeddings, representing words as distributed vectors. Notably, these word embeddings have consistently excelled in practical NLP tasks, profoundly shaping the field’s progress. Pre-trained language models (PLM) represent a subsequent phase in the evolution of language models following NLM. Early attempts at PLMs included ELMo [3], which was built on a Bidirectional LSTM architecture. However, with the advent of the transformer architecture [4], characterized by parallel self-attention mechanisms, the pre-training and fine-tuning learning paradigm has propelled PLM to prominence as the prevailing approach. These models are typically trained via self-supervision on extensive datasets, cementing their status as the primary methodology in the field.
@@
The Transformer architecture is exceptionally well-suited for scaling up models, and research analysis has revealed that increasing the model’s scale or training data size can significantly enhance its performance. Many studies have pushed the boundaries of model performance by continuously expanding the scale of PLM [5; 6; 7; 8]. As models grow larger, a remarkable phenomenon known as "emergence" occurs, wherein they exhibit astonishing performance [6]. These models are capable of generating high-quality text and possess robust learning and reasoning abilities. They can even tackle few-shot learning tasks through in-context learning (ICL) [6]. This remarkable capability enables their seamless application to a wide range of downstream tasks across diverse domains [9; 10; 11; 12].
@@
Pre-trained language models (PLMs) with significantly larger parameter sizes and extensive training data are typically denoted as Large Language Models (LLMs) [13; 14; 15]. The model size usually exceeds 10 billion (10B) parameters. A prominent milestone in the development of LLMs is exemplified by the GPT series [16; 5; 6; 17]. Notably, OpenAI released ChatGPT in November 2022, marking a pivotal moment in the era of LLMs and a game-changing moment in the field of artificial intelligence. ChatGPT has empowered current AI algorithms to achieve unprecedented levels of strength and effectiveness, reshaping the way humans employ or develop AI algorithms. Its emergence has captured the attention of the research community. However, owing to ChatGPT’s absence as an open-source platform, the principal way to use ChatGPT currently is by accessing it through OpenAI’s website at https://chat.openai.com or via their API interface. Training LLMs that can serve as alternatives to ChatGPT, or domain-specific LLMs, has become highly necessary [18; 19; 20; 21; 22; 1; 23; 24]. Training and deploying LLMs demand expertise in handling large-scale data and substantial practical experience in distributed parallel training [25; 26; 27]. This requirement emphasizes the need for researchers developing LLMs to possess significant engineering capabilities in addressing the challenges encountered during LLM development. Researchers who are interested in the field of LLMs must possess engineering skills or learn to collaborate effectively with engineers.
@@ 2.2 Prompt Learning
Prompt learning serves as a widely adopted machine learning approach, particularly in the field of NLP. At its core, this methodology involves guiding a model to produce specific behaviors or outputs through the careful design of prompt statements. It is commonly employed to fine-tune and guide pre-trained LLMs for executing particular tasks or generating desired results. Researchers have observed that the design of specific prompt statements can steer pre-trained models to perform various tasks, such as question-answering, text generation, and semantic understanding [37; 38; 39; 40; 41; 42; 43; 44; 45; 46; 47; 48]. The strength of this approach lies in its ability to adapt to different tasks through simple modifications to prompt statements, eliminating the need for retraining the entire model. For LLMs like the GPT series and other pre-trained models, prompt learning provides a straightforward and powerful means for model fine-tuning. By supplying appropriate prompts, researchers and practitioners can customize the model’s behavior, making it more suitable for specific domains or task requirements. In short, prompt learning is a machine learning approach that, builds upon pre-trained language models, and guides the model to perform various tasks through the design of prompt statements, offering increased flexibility for customizing model applications. In this Section, we will introduce the basic knowledge of prompt learning.
@@ 3.1.1 Dataset
Books: Two commonly utilized books datasets for LLMs training are BookCorpus [63] and Gutenberg [64]. These datasets include a wide range of literary genres, including novels, essays, poetry, history, science, philosophy, and more. Widely employed by numerous LLMs [7; 77], these datasets contribute to the models’ training by exposing them to a diverse array of textual genres and subject matter, fostering a more comprehensive understanding of language across various domains.
@@
Wikipedia: Wikipedia [73], a free and open online encyclopedia project, hosts a vast repository of high-quality encyclopedic content spanning a wide array of topics. The English version of Wikipedia is extensively utilized in the training of many LLMs [6; 7; 78], serving as a valuable resource for language understanding and generation tasks. Additionally, Wikipedia is available in multiple languages, providing diverse language versions that can be leveraged for training in multilingual environments.
@@ 3.1.2 Data preprocessing
Quality filtering: Filtering low-quality data is typically done using heuristic-based methods or classifier-based methods. Heuristic methods involve employing manually defined rules to eliminate low-quality data [82; 70]. For instance, rules could be set to retain only text containing digits, discard sentences composed entirely of uppercase letters, and remove files with a symbol and word ratio exceeding 0.1, and so forth. Classifier-based methods involve training a classifier on a high-quality dataset such as WebText [83] to filter out low-quality datasets.
@@ 3.2.2 Decoder-only Architecture
The Causal Decoder Architecture: In the Causal Decoder architecture, each token in the model input sequence can only attend to past input tokens and itself during the decoding process. It achieves unidirectional attention to the input sequence by using a specific mask as shown in Figure 1. In fact, different architectures are mainly implemented by configuring different mask matrices. The figure illustrates a comparison of mask configurations between the Encoder-decoder and Decoder-only architectures (including Casual Decoder and Prefix Decoder). The representative LLMs for the Causal Decoder architecture are the GPT series [16; 5; 6; 91; 17]. The GPT series of LLMs are currently known for their superior performance, with their foundational Causal Decoder architecture widely applied in other LLMs such as BLOOM [36], OPT [81], Gopher [82], and LLaMA [7].
@@ 3.3 Pre-training Tasks
In language modeling [16; 5; 6; 34], the model is required to predict the next word in a given context. This task enables the model to develop a nuanced understanding of language. Specifically, the model observes large amounts of textual data and attempts to predict the next word at each position in the text. This gradual learning process allows the model to capture the patterns and information inherent in language, encoding a vast amount of linguistic knowledge into its parameters. Once pre-training is complete, these model parameters can be fine-tuned for various natural language processing tasks to adapt to specific task requirements. The objective of language modeling is to train a model to maximize the likelihood of textual data. For a given text sequence, denoted as $w_{1},w_{2},...,w_{T}$, where $w_{t}$ represents the token at position $t$, $P(w_{t}|w_{1},w_{2},...,w_{t-1})$ is the probability of predicting $w_{t}$ given the preceding context $w_{1},w_{2},...,w_{t-1}$, the objective function for language modeling can be expressed using cross-entropy loss. Here, we define the objective as maximizing the conditional probability of the given text sequence:
@@
Language modeling serves as a prevalent pretraining objective for most LLMs. In addition to language modeling, there are other pretraining tasks within the realm of language modeling. For instance, some models [66; 35] use text with certain portions randomly replaced, and then employ autoregressive methods to recover the replaced tokens. The primary training approach involves the autoregressive recovery of the replaced intervals.
@@ 3.4.2 Mixed Precision Training
In recent years, to pre-train extremely large language models, some research [97] has begun to utilize 16-bit floating-point numbers (FP16) to reduce memory usage and communication overhead. FP16 has a smaller numerical range and lower precision in effective digits [98; 36], but computations tend to be faster than FP32. In general model training, FP32 is often used as the default representation for training parameters. However, in actual model training, the number of parameters in a model typically does not exceed the order of thousands, well within the numerical range of FP16. To improve computational speed, we can convert from FP32 to FP16. During parameter updates, the amount of the parameter is roughly equal to the gradient multiplied by the learning rate. The minimum value of FP16 is on the order of 1e-5. As the product of the gradient and learning rate is already well below the representation range of FP16, the parameter update would result in loss, known as underflow. Therefore, we represent the parameter update obtained by multiplying the gradient by the learning rate as FP32. We cannot directly add this high-precision parameter update to a lower-precision model, as this would still result in floating-point underflow. Consequently, we need to save an additional single-precision parameter on the optimizer. To accelerate both forward and backward passes in the model, half-precision parameters and gradients are used and passed to the optimizer for updating. The optimizer’s update quantity is saved as FP32, and we accumulate it effectively through a temporarily created FP32 parameter in the optimizer. After effective accumulation, it is then converted back to FP16 parameters.
@@ 3.5.2 Alignment Tuning
Due to LLMs being pre-trained on massive and diverse internet data, even though the training data undergoes some preprocessing, it is still challenging to guarantee the absence of biased or harmful content in terabyte-scale training datasets. Despite LLMs demonstrating impressive performance across various natural language processing tasks, they frequently exhibit behaviors diverging from human intent. This includes generating false information, producing expressions with bias or misleading content, and so on [91; 107]. To address these issues of LLMs displaying behaviors beyond human intent, alignment tuning becomes crucial [91; 108].
@@ 3.5.3 Parameter-efficient Tuning
Currently, large-scale PLMs such as ChatGPT [91; 17] continue to grow in scale. However, for the majority of researchers, conducting full fine-tuning on consumer-grade hardware has become cost-prohibitive and impractical. Unlike SFT and alignment tuning, the objective of parameter-efficient tuning is to reduce computational and memory overhead. This method involves fine-tuning only a small or additional subset of model parameters while keeping the majority of pre-trained parameters fixed, thereby significantly lowering computational and storage costs. It is noteworthy that state-of-the-art parameter-efficient tuning techniques have achieved performance levels comparable to full fine-tuning. Some common parameter-efficient tuning methods include Low-Rank Adaptation (LoRA) [110], Prefix Tuning [111] and P-Tuning [112; 113]. The adoption of these methods enables efficient model tuning even in resource-constrained environments, offering feasibility and efficiency for practical applications.
With the rise of LLMs, parameter-efficient tuning has garnered increasing attention, with LoRA being widely employed in the latest releases of LLMs. LoRA [110] and its related advancements [114; 115] are noteworthy and deserve attention.
@@ 3.6 Evaluation
Unlike in the past, large-scale deep learning models have a wider range of applications and stronger performance compared to ordinary models. However, with great power comes great responsibility, and evaluating these models has become more complex, requiring consideration of potential problems and risks from all aspects. Since the popularity of ChatGPT, many related studies have been published, including the survey and summary of LLMs evaluation in reference [117; 118], which is helpful for developing large-scale deep learning models. This section will introduce some testing datasets, evaluation directions and methods, and potential threats that need to be considered based on previous evaluation work on large models.
@@ 3.7 LLM Framework
Megatron-LM: Megatron-LM [94; 159; 160] is a deep learning library developed by NVIDIA for training large-scale language models.Megatron-LM presents their techniques including model and data parallelism, mixed-precision training, and FlashAttention for training very large transformer models. Specifically, it takes advantage of the structure of transformer networks to create a simple model parallel implementation by adding a few synchronization primitives and it enables training transformer models with billions of parameters and trains efficiently in PyTorch.It also performs an in-depth empirical analysis of their model and data parallel technique and demonstrates up to 76% scaling efficiency using 512 GPUs which can largely improve the training efficiency and speed, enabling efficient distributed training across GPUs.
In addition to the aforementioned frameworks, Colossal-AI [161] and FastMoE [162; 163] are also two popular frameworks for training LLMs. In principle, any deep learning framework that supports parallel computing can be used to train LLMs. Examples include PyTorch [164], TensorFlow [165; 166], PaddlePaddle [167], MXNet [168], OneFlow [169], MindSpore [170] and JAX [171].
@@ 4.3 Parallelism
Both inference and training can leverage parallelization techniques. Presently, parallelization techniques for inference primarily manifest across three dimensions: Data Parallelism, Tensor Parallelism, and Pipeline Parallelism. Data Parallelism primarily involves increasing the overall throughput of the inference system by adding more GPU devices [99; 95; 157; 183]. Tensor parallelism is a form of model parallelism where the model’s parameters are partitioned into multiple tensors, each computed on different processing units. This approach proves beneficial when dealing with models that are too large to fit into the memory of a single GPU. Tensor parallelism primarily involves increasing the number of devices horizontally through parallel computation to reduce latency [94]. Pipeline parallelism primarily involves vertically increasing the number of GPU devices through parallel computation to support larger models and enhance device utilization. Typically, it is combined with tensor parallelism to achieve optimal performance [96].

### v2

@@ 1 Introduction
Language modeling (LM) is a fundamental approach for achieving cognitive intelligence in the field of natural language processing (NLP), and its progress has been notable in recent years [1; 2; 3]. It assumes a central role in understanding, generating, and manipulating human language, serving as the cornerstone for a diverse range of NLP applications [4], including machine translation, chatbots, sentiment analysis, and text summarization. With the evolution of deep learning, the early statistical language models (SLM) have gradually transformed into neural language models (NLM) based on neural networks. This shift is characterized by the adoption of word embeddings, representing words as distributed vectors. Notably, these word embeddings have consistently excelled in practical NLP tasks, profoundly shaping the field’s progress. Pre-trained language models (PLM) represent a subsequent phase in the evolution of language models following NLM. Early attempts at PLMs included ELMo [5], which was built on a Bidirectional LSTM architecture. However, with the advent of the transformer architecture [6], characterized by parallel self-attention mechanisms, the pre-training and fine-tuning learning paradigm has propelled PLM to prominence as the prevailing approach. These models are typically trained via self-supervision on extensive datasets, cementing their status as the primary methodology in the field.
@@
The Transformer architecture is exceptionally well-suited for scaling up models, and research analysis has revealed that increasing the model’s scale or training data size can significantly enhance its performance. Many studies have pushed the boundaries of model performance by continuously expanding the scale of PLM [7; 8; 9; 10]. As models grow larger, a remarkable phenomenon known as "emergence" occurs, wherein they exhibit astonishing performance [8]. These models are capable of generating high-quality text and possess robust learning and reasoning abilities. They can even tackle few-shot learning tasks through in-context learning (ICL) [8]. This remarkable capability enables their seamless application to a wide range of downstream tasks across diverse domains [11; 12; 13; 14].
@@
Pre-trained language models (PLMs) with significantly larger parameter sizes and extensive training data are typically denoted as Large Language Models (LLMs) [15; 16; 17]. The model size usually exceeds 6-10 billion (6-10B) parameters. A prominent milestone in the development of LLMs is exemplified by the GPT series [18; 7; 8; 19]. Notably, OpenAI released ChatGPT in November 2022, marking a pivotal moment in the era of LLMs and a game-changing moment in the field of artificial intelligence. ChatGPT has empowered current AI algorithms to achieve unprecedented levels of strength and effectiveness, reshaping the way humans employ or develop AI algorithms. Its emergence has captured the attention of the research community. However, owing to ChatGPT’s absence as an open-source platform, the principal way to use ChatGPT currently is by accessing it through OpenAI’s website at https://chat.openai.com or via their API interface. Training LLMs that can serve as alternatives to ChatGPT, or domain-specific LLMs, has become highly necessary [20; 21; 22; 23; 24; 1; 25; 26]. Training and deploying LLMs demand expertise in handling large-scale data and substantial practical experience in distributed parallel training [27; 28; 29]. This requirement emphasizes the need for researchers developing LLMs to possess significant engineering capabilities in addressing the challenges encountered during LLM development. Researchers who are interested in the field of LLMs must possess engineering skills or learn to collaborate effectively with engineers.
@@ 2.2 Prompt Learning
Prompt learning serves as a widely adopted machine learning approach, particularly in the field of NLP. At its core, this methodology involves guiding a model to produce specific behaviors or outputs through the careful design of prompt statements. It is commonly employed to fine-tune and guide pre-trained LLMs for executing particular tasks or generating desired results. Researchers have observed that the design of specific prompt statements can steer pre-trained models to perform various tasks, such as question-answering, text generation, and semantic understanding [39; 40; 41; 42; 43; 44; 45; 46; 47; 48; 49; 50]. The strength of this approach lies in its ability to adapt to different tasks through simple modifications to prompt statements, eliminating the need for retraining the entire model. For LLMs like the GPT series and other pre-trained models, prompt learning provides a straightforward and powerful means for model fine-tuning. By supplying appropriate prompts, researchers and practitioners can customize the model’s behavior, making it more suitable for specific domains or task requirements. In short, prompt learning is a machine learning approach that, builds upon pre-trained language models, and guides the model to perform various tasks through the design of prompt statements, offering increased flexibility for customizing model applications. In this Section, we will introduce the basic knowledge of prompt learning.
@@ 3.1.1 Dataset
Books: Two commonly utilized books datasets for LLMs training are BookCorpus [65] and Gutenberg [66]. These datasets include a wide range of literary genres, including novels, essays, poetry, history, science, philosophy, and more. Widely employed by numerous LLMs [9; 79], these datasets contribute to the models’ training by exposing them to a diverse array of textual genres and subject matter, fostering a more comprehensive understanding of language across various domains.
@@
Wikipedia: Wikipedia [75], a free and open online encyclopedia project, hosts a vast repository of high-quality encyclopedic content spanning a wide array of topics. The English version of Wikipedia is extensively utilized in the training of many LLMs [8; 9; 80], serving as a valuable resource for language understanding and generation tasks. Additionally, Wikipedia is available in multiple languages, providing diverse language versions that can be leveraged for training in multilingual environments.
@@ 3.1.2 Data preprocessing
Quality filtering: Filtering low-quality data is typically done using heuristic-based methods or classifier-based methods. Heuristic methods involve employing manually defined rules to eliminate low-quality data [84; 72]. For instance, rules could be set to retain only text containing digits, discard sentences composed entirely of uppercase letters, and remove files with a symbol and word ratio exceeding 0.1, and so forth. Classifier-based methods involve training a classifier on a high-quality dataset such as WebText [85] to filter out low-quality datasets.
@@ 3.2.2 Decoder-only Architecture
The Causal Decoder Architecture: In the Causal Decoder architecture, each token in the model input sequence can only attend to past input tokens and itself during the decoding process. It achieves unidirectional attention to the input sequence by using a specific mask as shown in Figure 1. In fact, different architectures are mainly implemented by configuring different mask matrices. The figure illustrates a comparison of mask configurations between the Encoder-decoder and Decoder-only architectures (including Casual Decoder and Prefix Decoder). The representative LLMs for the Causal Decoder architecture are the GPT series [18; 7; 8; 93; 19]. The GPT series of LLMs are currently known for their superior performance, with their foundational Causal Decoder architecture widely applied in other LLMs such as BLOOM [38], OPT [83], Gopher [84], and LLaMA [9].
@@ 3.3 Pre-training Tasks
In language modeling [18; 7; 8; 36], the model is required to predict the next word in a given context. This task enables the model to develop a nuanced understanding of language. Specifically, the model observes large amounts of textual data and attempts to predict the next word at each position in the text. This gradual learning process allows the model to capture the patterns and information inherent in language, encoding a vast amount of linguistic knowledge into its parameters. Once pre-training is complete, these model parameters can be fine-tuned for various natural language processing tasks to adapt to specific task requirements. The objective of language modeling is to train a model to maximize the likelihood of textual data. For a given text sequence, denoted as $w_{1},w_{2},...,w_{T}$, where $w_{t}$ represents the token at position $t$, $P(w_{t}|w_{1},w_{2},...,w_{t-1})$ is the probability of predicting $w_{t}$ given the preceding context $w_{1},w_{2},...,w_{t-1}$, the objective function for language modeling can be expressed using cross-entropy loss. Here, we define the objective as maximizing the conditional probability of the given text sequence:
@@
Language modeling serves as a prevalent pretraining objective for most LLMs. In addition to language modeling, there are other pretraining tasks within the realm of language modeling. For instance, some models [68; 37] use text with certain portions randomly replaced, and then employ autoregressive methods to recover the replaced tokens. The primary training approach involves the autoregressive recovery of the replaced intervals.
@@ 3.4.2 Mixed Precision Training
In recent years, to pre-train extremely large language models, some research [99] has begun to utilize 16-bit floating-point numbers (FP16) to reduce memory usage and communication overhead. FP16 has a smaller numerical range and lower precision in effective digits [100; 38], but computations tend to be faster than FP32. In general model training, FP32 is often used as the default representation for training parameters. However, in actual model training, the number of parameters in a model typically does not exceed the order of thousands, well within the numerical range of FP16. To improve computational speed, we can convert from FP32 to FP16. During parameter updates, the amount of the parameter is roughly equal to the gradient multiplied by the learning rate. The minimum value of FP16 is on the order of 1e-5. As the product of the gradient and learning rate is already well below the representation range of FP16, the parameter update would result in loss, known as underflow. Therefore, we represent the parameter update obtained by multiplying the gradient by the learning rate as FP32. We cannot directly add this high-precision parameter update to a lower-precision model, as this would still result in floating-point underflow. Consequently, we need to save an additional single-precision parameter on the optimizer. To accelerate both forward and backward passes in the model, half-precision parameters and gradients are used and passed to the optimizer for updating. The optimizer’s update quantity is saved as FP32, and we accumulate it effectively through a temporarily created FP32 parameter in the optimizer. After effective accumulation, it is then converted back to FP16 parameters.
@@ 3.5.2 Alignment Tuning
Due to LLMs being pre-trained on massive and diverse internet data, even though the training data undergoes some preprocessing, it is still challenging to guarantee the absence of biased or harmful content in terabyte-scale training datasets. Despite LLMs demonstrating impressive performance across various natural language processing tasks, they frequently exhibit behaviors diverging from human intent. This includes generating false information, producing expressions with bias or misleading content, and so on [93; 109]. To address these issues of LLMs displaying behaviors beyond human intent, alignment tuning becomes crucial [93; 110].
@@ 3.5.3 Parameter-efficient Tuning
Currently, large-scale PLMs such as ChatGPT [93; 19] continue to grow in scale. However, for the majority of researchers, conducting full fine-tuning on consumer-grade hardware has become cost-prohibitive and impractical. Unlike SFT and alignment tuning, the objective of parameter-efficient tuning is to reduce computational and memory overhead. This method involves fine-tuning only a small or additional subset of model parameters while keeping the majority of pre-trained parameters fixed, thereby significantly lowering computational and storage costs. It is noteworthy that state-of-the-art parameter-efficient tuning techniques have achieved performance levels comparable to full fine-tuning. Some common parameter-efficient tuning methods include Low-Rank Adaptation (LoRA) [112], Prefix Tuning [113] and P-Tuning [114; 115]. The adoption of these methods enables efficient model tuning even in resource-constrained environments, offering feasibility and efficiency for practical applications.
With the rise of LLMs, parameter-efficient tuning has garnered increasing attention, with LoRA being widely employed in the latest releases of LLMs. LoRA [112] and its related advancements [116; 117] are noteworthy and deserve attention.
@@ 3.6 Evaluation
Unlike in the past, large-scale deep learning models have a wider range of applications and stronger performance compared to ordinary models. However, with great power comes great responsibility, and evaluating these models has become more complex, requiring consideration of potential problems and risks from all aspects. Since the popularity of ChatGPT, many related studies have been published, including the survey and summary of LLMs evaluation in reference [119; 120], which is helpful for developing large-scale deep learning models. This section will introduce some testing datasets, evaluation directions and methods, and potential threats that need to be considered based on previous evaluation work on large models.
@@ 3.7 LLM Framework
Megatron-LM: Megatron-LM [96; 161; 162] is a deep learning library developed by NVIDIA for training large-scale language models.Megatron-LM presents their techniques including model and data parallelism, mixed-precision training, and FlashAttention for training very large transformer models. Specifically, it takes advantage of the structure of transformer networks to create a simple model parallel implementation by adding a few synchronization primitives and it enables training transformer models with billions of parameters and trains efficiently in PyTorch.It also performs an in-depth empirical analysis of their model and data parallel technique and demonstrates up to 76% scaling efficiency using 512 GPUs which can largely improve the training efficiency and speed, enabling efficient distributed training across GPUs.
In addition to the aforementioned frameworks, Colossal-AI [163] and FastMoE [164; 165] are also two popular frameworks for training LLMs. In principle, any deep learning framework that supports parallel computing can be used to train LLMs. Examples include PyTorch [166], TensorFlow [167; 168], PaddlePaddle [169], MXNet [170], OneFlow [171], MindSpore [172] and JAX [173].
@@ 4.3 Parallelism
Both inference and training can leverage parallelization techniques. Presently, parallelization techniques for inference primarily manifest across three dimensions: Data Parallelism, Tensor Parallelism, and Pipeline Parallelism. Data Parallelism primarily involves increasing the overall throughput of the inference system by adding more GPU devices [101; 97; 159; 185]. Tensor parallelism is a form of model parallelism where the model’s parameters are partitioned into multiple tensors, each computed on different processing units. This approach proves beneficial when dealing with models that are too large to fit into the memory of a single GPU. Tensor parallelism primarily involves increasing the number of devices horizontally through parallel computation to reduce latency [96]. Pipeline parallelism primarily involves vertically increasing the number of GPU devices through parallel computation to support larger models and enhance device utilization. Typically, it is combined with tensor parallelism to achieve optimal performance [98].